# AdaSemiCD: An Adaptive Semi-Supervised Change Detection Method Based on Pseudo-Label Evaluation

### v1

@@ III-B Pseudo-label qualification metric
To enhance the effectiveness of pseudo-labels by accurately gauging their quality, a crucial step is the implementation of an evaluation metric. This metric will be instrumental in identifying reliable labels and determining the utility of each training sample. Unlike the scenario with labeled image pairs where real labels serve as a benchmark for assessing the model’s inference quality through average intersection over union, F1 scores, and so forth, pseudo-labels lack any such reference points and can only be compared to themselves. Therefore, we have devised a quantifiable evaluation metric specifically for pseudo-labels, designed to increase the total information entropy while taking into account factors like category imbalance and areas of confusion.
To evaluate the quality of pseudo-labels, a frequently employed technique for this is the computation of information entropy [57],
@@
where $P\left(x_{i}\right)$ is the output probability of a trained model on sample $x_{i}$. Our conviction is that a lower information entropy in the predicted value corresponds to a greater trustworthiness of the prediction outcome. On the contrary, a higher information entropy leads to a more varied prediction outcome, a more evenly distributed prediction probability on the same pixel, and a diminished distinction in the category.
@@
In our change detection task, the direct application of entropy often does not yield good results due to the significant challenge posed by category imbalance. As illustrated in Fig. 3, ten publicly available change detection datasets were evaluated, namely HRCUS-CD[58], LEVIR-CD[59], LEVIR-CD+[59],WHU-CD[60], EGY-CD[61], GZ-CD[26], CL-CD[62], CDD[63], SYSU-CD[64] and DSIFN-CD[65]. The ratios of the changed and unchanged categories are highly skewed. This significant category imbalance can lead to our model inadequately learning the target categories during training while disproportionately capturing feature distributions from background categories. As a result, there is an increased tendency to classify pixels as belonging to the background class during inference. Consequently, although the overall accuracy (OA) of the model appears high, the performance on the Intersection over Union (IoU) across various categories remains unsatisfactory. To minimize the influence of this category imbalance when assessing the quality of pseudo-labels, we assigned different weights to the two categories when computing information entropy as Eq. 7.
@@
where $w_{0}$ and $w_{1}$ represent the proportions of pixels in the current batch that belong to the unchanged and changed categories, respectively. The formulas for their calculation are given by Eq. 8 and Eq. 9.
@@
Moreover, pixels from diverse areas contribute in varying manners. Pixels of greater importance, like those in boundary, target, and regions similar to the background, who are offenly predicted with large uncertainty, should receive more emphasis. Initially, we enhance the impact of these essential regions by computing the absolute value of the disparity between the prediction probabilities of the two categories.
@@
Here, $abs$ signifies absolute value, a measure taken to prevent the inconsistencies that may arise from varying changes both prior to and following a phase. Then doing pixel-level product operation with information entropy, we get the uncertainty map of image $(x_{i})$,
@@
Clearly, in pixel locations with low information entropy, this process won’t result in notable alterations, whereas in pixel locations with high information entropy, the absolute value of the predictive probability post-subtraction will be significantly reduced, and the immediate amalgamation will diminish its importance. Consequently, we adopt the inverse of rectified information entropy as its depiction.
To sum up, with the goal of evaluating the quality of pseudo-labels for change detection, we put forward a measurable calculation metric $U$ to gauge uncertainty, which enhances the overall information entropy considering factors like category imbalance and focus on areas of confusion.
@@ III-C AdaFusion: Adaptive sample fusion
Image fusion is frequently employed to augment samples and boost generalization, with CutMix[34] and MixUp[66] being typical examples. In this study, our aim is to leverage image fusion techniques to exclude the unreliable areas in a training sample. This procedure involves two steps: region selection, which decides the location of the operation, and image selection, which decides the content to be used.
Adaptive selection of fusion region. Contrary to the conventional CutMix technique, which arbitrarily chooses blending regions, our approach is more refined. Initially, we would select a bounding box with random size, then slide the window, and choose the area where the total of uncertainty is highest as the region to be combined, using Eq. 11.
These regions are frequently boundaries or intricate areas where it’s challenging to identify targets. The CutMix fusion can guarantee the variety of samples and simultaneously decrease the unsupervised noise.
Adaptive selection of fusion contents. For the region of maximum uncertainty, we have the option to choose a substitute from its corresponding samples from the label section $B_{l}$ or the unlabeled section with higher probabilities $B_{u}$. This strategy prevents overuse of only labeled samples and further mitigates the risk of overfitting. The focus of the content here is confined to the contrast between labeled and unlabeled image pairs, rather than a more detailed distinction (such as the level of the image or the level of the object). This is because we suggest that too much human intervention could potentially undermine the model’s capacity to generalize. The choice of fusion content is mainly accomplished by adaptively adjusting a threshold that determines whether to fuse with a labeled image pair. We directly use the computed uncertainty as the threshold, and if the randomly generated probability exceeds the total uncertainty of the sample, the labeled images in the training batch are randomly selected as the fusion content. If not, other unlabeled image pairs in the batch are randomly selected. It is easy to understand that the higher the sample’s reliability, the better the quality of the pseudo-label. We consider it reliable enough to avoid the necessity for fusion with the limited labeled images, which could potentially increase the likelihood of overfitting. Samples with higher uncertainty come with more noise, and the addition of new content can significantly reduce the noise density.
@@ III-D AdaEMA: Adaptive EMA update strategy
Within the teacher-student framework, the teacher model is obtained by integrating the moving exponential average of the student model across the time series. This approach is more stable and reliable compared to dual models with shared weights. However, our expectation for Mean-Teacher is to reach the optimal state of coevolution, where the teacher model is a cumulative representation of the evolving student model. The crucial element in achieving this co-evolution is whether the student model progresses or regresses each time the teacher model gets updated. So, how can we assess the state of the student model? This question is inherently related to the validation phase of the training process. Typically, after training for several epochs, we evaluate the current model, feed samples from the validation set into the model for inference, and compare the outcomes with the actual labels to determine accuracy. However, if we perform such validation after every iteration, it becomes highly computationally intensive, leading to a significant increase in training time. Lowering the number of samples in the validation set (to just a few pairs) might address the feasibility issue, but it also presents new challenges when the sample size is too small to thoroughly assess the model’s performance. So, is there a middle ground that can merge these two concepts?
@@
0:    Student model $M_{stu}^{\theta}$, Teacher model $M_{tea}^{\theta^{\prime}}$ The set of training samples for current batch =$\mathcal{B}=\left\{\mathcal{B}_{l},\mathcal{B}_{u}\right\}$;
0:    Updated Teacher model, $M_{tea}^{\hat{\theta}}$;
1:  Calculate the loss $\mathcal{L}_{s}$ on labeled samples $\mathcal{B}_{l}$ with Eq. 1
2:  Calculate the loss $\mathcal{L}_{u}$ on unlabeled samples $\mathcal{B}_{u}$ with Eq. 2
3:  Update the student model $M_{stu}^{\theta}$ to $M_{stu}^{\theta^{\prime}}$ by SGD to minimize the total loss following Eq. 3
4:  Calculate the uncertainty ${U}_{tea}$ of pseudo-labels generated on $\mathcal{B}_{u}$ by the teacher model $M_{tea}^{\theta^{\prime}}$according to Eq.11
5:  Calculate the uncertainty ${U}_{stu}$ on $\mathcal{B}_{u}$ by updated student model $M_{stu}^{\theta^{\prime}}$ according to Eq.11
6:  Calculate the upper bound on the probability of update $\tau$ according to Eq. 12 and Eq. 13;
@@
8:     Update the teacher model to $M_{tea}^{\hat{\theta}}$ by Eq. 5
@@
10:     $M_{tea}^{\hat{\theta}}=M_{tea}^{\theta^{\prime}}$
@@
12:  return  $M_{tea}^{\hat{\theta}}$.
At each training phase, we initially update the student model’s parameters, denoted as $M_{stu}^{\theta}$, following the training strategy described in section III-A, resulting in $M_{stu}^{\theta^{\prime}}$. Next, we evaluate both the updated student model $M_{stu}^{\theta^{\prime}}$ and the teacher model $M_{tea}^{\theta^{\prime}}$ on the current batch of unlabeled training samples $\mathcal{B}_{u}$. We determine the ensuing uncertainty map using Eq. 11, and denote these as $U_{tea}$ and $U_{stu}$. The progression is marked by changes in uncertainty, as illustrated in Eq. 12.
$\varepsilon=\frac{\sum{U_{stu}}-\sum{U_{tea}}}{|B_{u}|}.$ (12)
If $\varepsilon$ is below zero, the model either degrades or oscillates. In contrast, when $\varepsilon$ is zero or higher, the model progresses. Ultimately, only the student models that have evolved successfully are selected to refresh the teacher model, with the maximum probability of update $\tau$ being derived as presented in Eq. 13.
Here, $\epsilon=1e-5$ is used to prevent the divisor from being zero, and $iter$ denotes the present iteration count. If the model parameters are not selected for updating the Teacher model, they are not discarded outright. Instead, they are given a lower probability of being updated. With this upper bound regulation (ranging from 0 to 1), some randomness is introduced to determine the likelihood of updating the teacher network’s parameters. Further details are available in Algorithm 1.
IV Experiement
This section pertains to the experimental validation we performed. Initially, we present the assessment standards for experimental outcomes and the experimental datasets. Subsequently, we executed ablation experiments to validate the significance of each component in the methodology. Ultimately, we draw comparisons with the cutting-edge semi-supervised change detection techniques, using measurable experimental metrics and perceptible visual outcomes.
@@ IV-A1 Datasets
Our method is empirically tested on ten benchmark datasets, namely the LEVIR-CD [59], LEVIR-CD+ [59], WHU-CD[60], EGY-CD[61], HRCUS-CD[58], Change Detection Dataset(CDD)[63], GZ-CD[26], DSIFN-CD[65], SYSU-CD[64] and CL-CD[62]. Among them, these datasets cover different resolutions (0.03m-2.0m), different data sizes (2400-20000 pairs), different annotation categories (binary or multiclass), and different time spans between image pairs (1-16 years), as summarized in the Table I.
@@
Building HRCUS-CD [58] 0.5m 256 $\times$ 256 11388 Varing Link
@@ IV-A2 Evaluation Metrics
To facilitate a straightforward comparison with the latest cutting-edge techniques, we employed intersection over union (IoU) scores and overall accuracy (OA) on evaluation datasets. Given the severe imbalance in CD categories and our primary concentration on the changed area, we use the IoU of the change category $IoU^{c}$, the calculation formula are as follows:
$\mathrm{IoU}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}+\mathrm{FN}}$ (14)
$\mathrm{OA}=\frac{\mathrm{TP}+\mathrm{TN}}{\mathrm{TP}+\mathrm{FP}+\mathrm{TN}+\mathrm{FN}}$ (15)
Where $\mathrm{TP}$ represents the positive sample correctly predicted (the correct changing pixel), $\mathrm{TN}$ refers to the negative sample correctly predicted (the correct unchanged pixel), and $\mathrm{FP}$ denotes the positive sample wrongly predicted (the unchanged pixel wrongly detected), $\mathrm{FN}$ represents the negative sample wrongly predicted (the pixel missed as the unchanged pixel). For both metrics, the larger the value, the better the change detection performance of the model. In the change detection task, OA is high overall due to binary classification and unbalanced category envy.
@@ IV-A3 Implementation Details
The common components of this experiment were tested under the same conditions as the methods we intend to compare them with. To prevent network structure from causing interference, we utilized ResNet50+PPM as our change detection network follow [25] and [27]. We set the initial learning rate at 0.01, which linearly decreases to 1e-4 with a momentum of 0.9, and trained all methods using the SGD optimizer. All models underwent 80 epochs of training, with each batch having a mini-batch size of 8 for both labeled and unlabeled training sets. Furthermore, all the enhancements employed were identical to those used in FPA[27], which include random flip, random resize ranging from 0.5 to 2.0, and random crop among the weak augmentations, along with the nine strong augmentations cited in [67]. Given the limited amount of labeled data, the model is highly sensitive to the pseudo-labels’ filtering threshold, leading us to set 0.95 as the threshold for all models in the experiments conducted on the ten datasets and we follow the experience in [25] to set $\phi$ to 5. All our experiments were carried out using PyTorch on four NVIDIA GeForce RTX 3090 GPUs.
@@ IV-B Compare with the SOTA methods
In order to prove the superiority of our proposed method, we compare it with several of the most advanced semi-supervised change detection methods. [26], [25] and [27] refresh the last semi-supervised CD performance in chronological order. In addition, we also compare with two semi-supervised semantic segmentation modding methods[37] and [68].
@@ IV-B1 Quantitative results
We conducted a complete experiment on ten datasets, Table II and Table III report our experimental results on the Building CD datasets and mutilclass CD datasets respectively. Sup.Only is the result of supervised training only on a limited subset of labeled dataset, while Oracle is the result of fully supervised training on the entire training dataset. It is worth noting that our method achieves the state-of-the-art performance in almost all partition settings across the all datasets. To be sure, all the semi-supervised CD methods performed better than the supervised methods in the corresponding partition setting in most cases, which proves the effectiveness of the semi-supervised methods in learning from a large number of unlabeled training samples, at the same time, the lead achieved by our method proves that our adaptive design for how to learn from unlabeled samples to more correct guidance is of practical significance.
@@
In addition, from the experimental results, the overall accuracy(OA) of all methods on these binary change detection datasets is very high, because the background class accounts for the majority. OA is often directly corresponding to $IoU_{c}$, and a slight improvement in OA is reflected in a relatively significant improvement in $IoU_{c}$. Our AdaSemiCD improves OA in almost all datasets because it greatly eliminates noise and the guidance of false signals.
@@
Oracle $IoU^{c}$=77.9 and OA=98.77
Multiclass CD Datasets: As shown in TableIII, our AdaSemiCD achieved SOTA performance on almost all datasets and all settings, just the improvement was less than on the binary building change detection datasets. This is due to the complexity of the task itself, and one of the most obvious is that in some cases, some of the previous semi-supervised change detection methods failed and performed worse than the baseline under full supervision. However, our semi-supervised CD approach can still leverage additional unlabeled samples to improve performance.
Another point worth noting is that in these multiclass change detection datasets, the overall accuracy of all methods is reduced, and OA and $IoU_{c}$ are no longer directly corresponding. One of the phenomena is that $IoU_{c}$ reaches the highest, but OA does not, which is due to its different emphasis. OA pays more attention to the global accuracy, that is, the detection of background categories. $IoU_{c}$ is more concerned with the detection of the target class, which is more important in CD than the background class. As you can see in DSIFN-CD, our approach has achieved good results on the $IoU_{c}$.

### v2

@@ III-B Pseudo-label Qualification Metric
To enhance the effectiveness of pseudo-labels by accurately gauging their quality, a crucial step is the implementation of an evaluation metric. This metric plays a key role in identifying reliable labels and determining the value of each training sample. Unlike labeled image pairs, where true labels serve as a benchmark for assessing the model’s performance, pseudo-labels lack such reference points and can only be compared to themselves. To evaluate the quality of pseudo-labels, a common technique is the computation of information entropy [51],
@@
where $P\left(x_{i}\right)$ is the output probability of a trained model on sample $x_{i}$. We propose that reduced information entropy in the predicted values implies greater reliability of the prediction. In contrast, increased information entropy points to a prediction with more uncertainty, indicating a more balanced probability distribution across the pixels and resulting in decreased confidence in the model’s predictions.
@@
In the CD task, directly applying entropy may not yield the optimal results due to the significant challenge posed by class imbalance, as illustrated in Fig. 2. The ratios of changed to unchanged categories are highly skewed. This imbalance may lead our model to learn the target categories inadequately during training while disproportionately capturing feature distributions from the background categories. As a result, the model tends to classify pixels as background during inference. To reduce the influence of this class imbalance when assessing the quality of pseudo-labels, we assign different weights to the two categories when computing information entropy,
@@
where $w_{0}$ and $w_{1}$ represent the proportions of pixels in the current batch that belong to the unchanged and changed categories, respectively. The formulas for their calculation are:
@@
Furthermore, pixels from different regions play distinct roles. Pixels that are more crucial, like those found on edges, targets, or areas resembling the background—frequently predicted with higher uncertainty—should be prioritized. To achieve this, we start by enhancing the significance of these key regions by calculating the absolute difference between the prediction probabilities of the two classes,
@@
In this context, $abs$ represents the absolute value function, employed to prevent inconsistencies that could occur due to differing changes before and after a phase. By carrying out a pixel-wise multiplication with the information entropy, we derive the uncertainty map for the image $x_{i}$,
@@
In areas characterized by low information entropy, this procedure is unlikely to cause significant changes. However, in regions where information entropy is high, the operation enhances the effect.
To summarize, aiming to assess the quality of pseudo-labels in change detection, we propose a quantifiable metric $U$ to evaluate uncertainty. This metric improves overall information entropy by considering elements such as class imbalance and regions of confusion.
@@ III-C AdaFusion: Adaptive Sample Fusion
Image fusion is widely used to augment samples and enhance generalization, with CutMix[31] and MixUp[52] being typical examples. In this study, our aim is to apply image fusion techniques to exclude unreliable areas from the training samples. This process consists of two steps: region selection, which determines the location for the operation, and image selection, which specifies the content to be used.
Adaptive selection of fusion region. Contrary to the conventional CutMix technique, which arbitrarily selects blending regions, our approach is more refined. First, we initialize a bounding box of random size, then slide the window to identify the area with the highest total uncertainty using Eq. (12), as the region to be fused. These regions typically include boundaries or complex areas where target identification is difficult. The selection approach guarantees a varied sample and, at the same time, decreases unsupervised noise.
Adaptive selection of fusion contents. For the region of maximum uncertainty, we can select a substitute from either the corresponding samples in the labeled set $B_{l}$ or the unlabeled set $B_{u}$ with higher reliability. This strategy helps prevent overreliance on labeled samples and further mitigates the risk of overfitting. The selection of fusion content is guided by an adaptive threshold that determines whether to fuse with a labeled image pair. We directly use the computed uncertainty as the threshold, and if the randomly generated probability exceeds the total uncertainty of the sample, labeled images in the training batch are selected as the fusion content. Otherwise, other unlabeled image pairs of higher quality are chosen. It is clear that the higher the reliability of a sample, the better the quality of the pseudo-label. If a sample is considered reliable enough, we avoid fusing it with limited labeled images, which could increase the risk of overfitting. Samples with higher uncertainty contain more noise, and incorporating new content is expected to reduce the noise effectively.
@@ III-D AdaEMA: Adaptive EMA Update Strategy
Within the MT framework, the teacher model’s update process involves incorporating the student’s exponential moving average over time. Our goal is to reach a state of optimal coevolution, where the teacher acts as an aggregate reflection of the student model’s progress. A pivotal factor in this coevolution is whether the student model advances or declines with each update of the teacher. Evaluating the student model’s state is closely tied to the training process’s validation phase. Typically, after several training epochs, we evaluate the model by using samples from the validation set, performing inference, and comparing the output against true labels to determine accuracy. Conducting this validation after every iteration incurs a significant computational cost and extends training duration. One may address this by reducing the validation set size to just a few pairs; however, this risks insufficiently assessing model performance if the sample is too small. The challenge lies in striking a balance between these considerations.
During each training phase, we commence by updating the student model, denoted as $M_{stu}^{\theta}$, according to the strategy outlined in section III-A, which results in the updated model $M_{stu}^{\theta^{\prime}}$. Subsequently, we assess both the modified student model, $M_{stu}^{\theta^{\prime}}$, and the teacher model, $M_{tea}^{\theta}$ on the current set of unlabeled training samples, $\mathcal{B}_{u}$. Using Eq. (12), we calculate the corresponding uncertainty maps, referred to as $U_{tea}$ and $U_{stu}$. The changes in the student model’s development are captured through fluctuations in uncertainty,
$\varepsilon=\frac{\sum{U_{stu}}-\sum{U_{tea}}}{|B_{u}|}.$ (13)
Additionally, we introduce a probability $\tau$ to regulate model updates according to the reliability of the student model, defined as
Here, $iter$ denotes the current iteration count and $\epsilon=1e-5$. When $\varepsilon$ is greater than zero, the student model progresses, allowing a straightforward update of $M_{tea}^{\theta}$. In contrast, if $\varepsilon$ is zero or less, the student model encounters either regression or fluctuation. To gauge the probability of altering $M_{tea}^{\theta}$, we incorporate some randomness. The adaptive updating details are provided in Algorithm 1.
@@
0:    Student model $M_{stu}^{\theta}$, Teacher model $M_{tea}^{\theta}$ The set of training samples for the current batch, =$\mathcal{B}=\left\{\mathcal{B}_{l},\mathcal{B}_{u}\right\}$
0:    Updated Teacher model, $M_{tea}^{\theta^{\prime}}$;
1:  Calculate the supervised loss $\mathcal{L}_{s}$ on labeled samples $\mathcal{B}_{l}$ using Eq. (1);
2:  Calculate the unsupervised loss $\mathcal{L}_{u}$ on unlabeled samples $\mathcal{B}_{u}$ using Eq. (2);
3:  Update the student model $M_{stu}^{\theta}$ to $M_{stu}^{\theta^{\prime}}$ using SGD to minimize the total loss, as described in Eq. (3).
4:  Calculate the uncertainty ${U}_{tea}$ of the pseudo-labels generated on $\mathcal{B}_{u}$ by the teacher model $M_{tea}^{\theta}$ as described in Eq. (12);
5:  Calculate the uncertainty ${U}_{stu}$ of the pseudo-labels generated on $\mathcal{B}_{u}$ by updated student model $M_{stu}^{\theta^{\prime}}$ as described in Eq. (12);
6:  Calculate the upper bound of the update probability $\tau$ according to Eq. (13) and Eq. (14);
@@
8:     Update the teacher model to $M_{tea}^{\theta^{\prime}}$ by EMA;
@@
10:     $M_{tea}^{\theta^{\prime}}=M_{tea}^{\theta}$
@@
12:  return  $M_{tea}^{\theta^{\prime}}$.
IV Experiment
@@ IV-A1 Datasets
Figure 3: Typical examples of ten change detection datasets we use, with Pre-event image, Post-event image, and Ground Truth.
Our method is empirically tested on ten benchmark datasets, namely the LEVIR-CD [53], LEVIR-CD+ [53], WHU-CD[54], EGY-CD[55], HRCUS-CD[56], CDD [57], GZ-CD[22], DSIFN-CD[58], SYSU-CD[59], and CL-CD[60]. As summarized in Table I, these datasets cover different resolutions (0.03m-2.0m), different data sizes (1-20000 pairs), different annotation categories (binary or multiclass), and different time spans (2-16 years). Fig. 3 shows some classic sample images for each dataset.
@@
Binary HRCUS-CD [56] 0.5m 256 $\times$ 256 11388 Varing Link
@@ IV-A2 Evaluation Metrics
For easy comparison with the most advanced techniques, we utilized overall accuracy (OA) to assess general performance. Due to the significant imbalance in CD categories and our main focus on the altered area, we applied the intersection over union for the changed category $IoU^{c}$. The calculation formulas are provided below:
$IoU^{c}=TP/(TP+FP+FN)$ (15)
$OA=(TP+TN)/(TP+FP+FN+TN)$ (16)
Where $\mathrm{TP}$ represents the positive sample correctly predicted (the correct changing pixel), $\mathrm{TN}$ refers to the negative sample correctly predicted (the correct unchanged pixel), and $\mathrm{FP}$ denotes the positive sample wrongly predicted (the unchanged pixel wrongly detected), $\mathrm{FN}$ represents the negative sample wrongly predicted (the pixel missed as the unchanged pixel). For both metrics, the larger the value, the better the change detection performance of the model.
@@ IV-A3 Implementation Details
We employ ResNet50+PPM as our change detection framework, as referenced in [21] and [23]. The learning rate starts at 0.01 and decreases linearly to 1e-4, with momentum kept at 0.9. Training of all competing approaches is conducted using the SGD optimizer over 80 epochs. For both labeled and unlabeled data, the mini-batch size is set at 8. Moreover, the augmentations applied are the same as those in FPA [23], which consist of weak augmentations such as random flip, random resizing from 0.5 to 2.0, and random cropping, along with nine robust augmentations [61]. A pseudo label threshold of 0.95 is adopted for all models in our studies. When calculating loss, $\phi$ is fixed at 5. The entire set of experiments was executed using PyTorch on four NVIDIA GeForce RTX 3090 GPUs.
@@ IV-B Compare with the SOTA Methods
To demonstrate the advantages of our proposed method, we have conducted a comparison with a number of leading semi-supervised change detection techniques [22, 21, 23, 35, 62].
@@ IV-B1 Quantitative Results
Tables II and III display our experimental outcomes for both the Building CD datasets and the multi-class CD datasets. The ‘Sup. only’ results denote supervised training outcomes from a restricted portion of the labeled dataset, whereas ‘Oracle’ represents full supervision results using the entire training dataset. It’s notable that our approach achieves SOTA performance in nearly every partition configuration across these datasets. Remarkably, in most scenarios, all semi-supervised CD approaches outperformed the supervised ones within the same partition setting, confirming the efficacy of semi-supervised methods in leveraging numerous unlabeled training samples. Additionally, the superiority of our method highlights the practical importance of our adaptive strategy for learning more effectively from unlabeled data.
Building CD Datasets: As presented in Table II, the proposed AdaSemiCD framework demonstrates outstanding performance across nearly all building CD datasets. Notably, on LEVIR-CD, LEVIR-CD+, and WHU-CD datasets, AdaSemiCD achieves a significant improvement in $IoU^{c}$, with gains of 3.1, 1.3, and 1.7 percentage points, respectively. On EGY-CD and HRCUS-CD datasets, AdaSemiCD maintains an overall leading position with average improvements of 0.75 and 0.4 points, respectively. However, performance on certain configurations is marginally lower than SOTA methods. This limitation can be attributed to the unique challenges posed by these datasets: the EGY-CD dataset suffers from overexposure, which often blends buildings with the background, increasing the difficulty of accurate discrimination. Similarly, the HRCUS-CD dataset is affected by vegetation occluding the changing areas of some buildings. These challenges hinder the model’s ability to generate reliable pseudo-labels when trained with a limited number of real labels.
@@
Furthermore, as indicated by the experimental results, the overall accuracy of all methods on these binary change detection datasets is notably high, primarily due to the dominance of the background class. OA often correlates closely with $IoU^{c}$, where even a slight improvement in OA translates into a relatively significant improvement in $IoU^{c}$. Our AdaSemiCD consistently enhances OA across nearly all datasets, as it effectively reduces noise and minimizes the influence of false signals, leading to more accurate predictions.
@@
Oracle $IoU^{c}$=50.1 and OA=95.66
Multiclass CD Datasets: As shown in Table III, our proposed AdaSemiCD demonstrates optimal performance in the majority of scenarios, despite the increased complexity of multi-category change detection tasks compared to binary building change detection tasks. Notably, optimal results were achieved across all experimental settings on the CDD dataset, with $IoU^{c}$ and OA increasing by an average of 1.8 percentage points and 0.25 percentage points respectively, which can be attributed to its exceptionally accurate annotations, as illustrated in Fig. 3. This high-quality labeling is particularly advantageous for pseudo-label-based semi-supervised methods. Furthermore, AdaSemiCD exhibits overall performance improvements on the SYSU-CD and CL-CD datasets. However, its performance is slightly below the SOTA results in certain experimental configurations. This can be attributed to the inherent challenges in identifying some mountain and land changes present in these datasets. These changes are often large-scale, and misclassification in such regions may result in fluctuations in performance metrics.
An additional significant observation is that, when the overall accuracy declines, the correlation between OA and $IoU^{c}$ becomes less straightforward. In particular, there are instances where $IoU^{c}$ reaches its maximum value, while OA does not, underscoring their different focuses. OA is mainly concerned with overall accuracy, highlighting the detection of background categories, whereas $IoU^{c}$ prioritizes the precise detection of target classes, which is more vital for change detection tasks than identifying background categories. This difference is apparent in the DSIFN-CD dataset, where our method demonstrates strong performance on the $IoU^{c}$ metric.
[…]